A triangular shaped stack of tin cans has 8 cans in the first row and 8 rows in all. In each successive row one can is removed.

What is the explicit rule for this situation, and how many cans will be in the5th

row?

Drag and drop the answers into the boxes to match the situation.

Explicit rule

Number of cans in the 5th

 row

a

n

=9−n

a

n

=8−n

a

n

=8−8n

a

n

=9−9n

345





1
Expert's answer
2021-05-11T09:02:05-0400

A triangular shaped stack of tin cans has 8 cans in the first row and 8 rows in all.


In each successive row one can is removed.

i.e. in first row there are 8 cans

in second row there are 7 cans

in third row there are 6 cans


The explicit formula which tells about the number of cans in row n is:

"a_n=9-n"

(Since "a_1=" 8 is satisfied only by this formula)

Number of cans in fifth row are:

"a_5=9-5=4"


Hence, number of cans in fifth row equals:

4